Solution of Linear Systems by the Echelon Method.
4x-y+3z=-2
3x+5y-z=15
-2x+y+4z=14
And:
x+2y+3z=11
2x-y+z=2
Thank you in advance!

Answer by Alan3354(69443)

Solution of Linear Systems by the Echelon Method.
4x-y+3z=-2
3x+5y-z=15
-2x+y+4z=14
---------------
You can check your answers:
x = -1
y = 4
z = 2
-----------
And:
x+2y+3z=11
2x-y+z=2
---------
You need a 3rd eqn